OpenStudy (anonymous):

Find the product of (x + 3)(x − 3). x2 − 6x + 9 x2 + 6x + 9 x2 + 9 x2 − 9

OpenStudy (hoblos):

use difference of two squares formula (a-b)(a+b) = a^2 - b^2 so (x + 3)(x − 3) = x2 − 9